# AG-SLAM: Active Gaussian Splatting SLAM

### v1

@@ Abstract
We present AG-SLAM, the first active SLAM system utilizing 3D Gaussian Splatting (3DGS) for online scene reconstruction. In recent years, radiance field scene representations, including 3DGS have been widely used in SLAM and exploration, but actively planning trajectories for robotic exploration is still unvisited. In particular, many exploration methods assume precise localization and thus do not mitigate the significant risk of constructing a trajectory, which is difficult for a SLAM system to operate on. This can cause camera tracking failure and lead to failures in real-world robotic applications. Our method leverages Fisher Information to balance the dual objectives of maximizing the information gain for the environment while minimizing the cost of localization errors. Experiments conducted on the Gibson and Habitat-Matterport 3D datasets demonstrate state-of-the-art results of the proposed method.
@@ 1 Introduction
Being able to autonomously explore and map an environment while localizing within that map is a core skill for a mobile robot. This problem is known as active Simultaneous Localization and Mapping (active SLAM). Active SLAM lies at the intersection of exploration and SLAM, although it introduces challenges that are not present in either – that is, a trade-off between exploration and reducing the uncertainty of the estimated state which includes the agent’s pose and the map of the environment. Classical active SLAM methods often define objectives to reduce the uncertainty of the state estimate based on its covariance [10; 40; 3; 55; 5]. The covariance is readily available for classical SLAM systems which use filters to update the state estimate. However, many recent visual SLAM systems instead use a non-linear rendering loss to update the state estimate, which makes the covariance difficult to obtain. In particular, systems of this type using 3D Gaussian Splatting (3DGS) [21] for the scene representation have been developed [30; 59; 20; 16] which allow for high-fidelity rendering of novel views of the scene. In addition to cases where the reconstruction of a scene is desired for its own sake, a 3DGS scene representation extended to support open-vocabulary semantic segmentation [67; 46; 38; 63] and can be used as a basis for language-specified robotics tasks, for example 3DGS have already been used for mobile manipulation [14]. Many existing methods for such tasks currently rely on pre-scanning the scene [42; 26; 15], so the ability to efficiently and autonomously create a 3DGS representation of the scene can support work using 3DGS for these tasks. While an existing active SLAM algorithm could be used to construct a 3DGS representation using the actions and estimated poses as inputs this is less efficient and, we argue, less effective than an active SLAM system which is specifically designed for a 3DGS scene representation. We thus present the first 3DGS-based active SLAM system, allowing us to autonomously create a scene representation of a novel environment from which we can render high-fidelity color and depth images.
Classical approaches such as frontier-based exploration and A* algorithms are still used in active SLAM systems for their efficiency and simplicity [47; 2; 53; 49; 22]. However, algorithms that use only simple heuristics cannot determine the information gain, limiting their objective to simply improving coverage or having minimal travel distances. We propose to formulate this problem as an active learning problem, use heuristic approaches to efficiently generate a large number of feasible paths as candidates, and employ an uncertainty-aware algorithm to determine the best path for both localization and mapping. This allows us to balance the dual objectives of exploration and localization uncertainty reduction by using frontiers and expected information gain to drive exploration and our novel path selection algorithm to minimize the uncertainty of the state estimate. There have been previous methods for quantifying the uncertainty of radiance fields for reconstructing scenes from given data [44; 45; 50; 12], active view selection [34; 19; 50; 25] and active reconstruction [60] or mapping [61] of scenes with given localization, and for active SLAM of small scenes with an inward-facing camera [64]. However, all the prior methods only model the uncertainty of the scene representation, whereas we also model the localization uncertainty. In addition, we consider uncertainty over paths not only single views.
To validate our approach, we evaluate our method on scenes from the Gibson [57] and Habitat-Matterport 3D [39] dataset quantitatively and qualitatively. We show superior reconstruction quality in various metrics compared to several baselines and recent state-of-the-art methods. In particular, we compare to Active Neural SLAM [4], ExplORB [36], UPEN [11], active-INR [61] and frontier-based exploration [58] using the ratio of the frontier area to the distance to it as the selection criteria. In all cases, to make a fair comparison of the rendering quality, we only use the method to select actions and keep the SLAM backend, which is used for the final rendering evaluation, the same. Our contributions can be summarized as follows:
  • We propose AG-SLAM, an active SLAM system that uses a 3D Gaussian representation. To the best of our knowledge, we are the first to study active SLAM problems with a 3D Gaussian representation.
  • We derive an objective function for paths in our 3D Gaussian representation that effectively balances the information gain for exploration and the cost of possible localization errors.
@@ 2 Related work
Visual SLAM
[…]
The success of radiance fields in novel view synthesis has made them a popular scene representation for dense SLAM. The first SLAM system using a radiance field as the only scene representation was iMAP [48], using a multi-layer perceptron (MLP) as its scene representation. NICE-SLAM [68] used an explicit hierarchical feature grid along with pre-trained MLPs to represent the scene, allowing it to better represent large-scale scenes. NICE-SLAM [68] was extended to only require RGB data instead of RGB-D in NICER-SLAM [69] by making use of monocular depth and normal estimators and optical flow loss. GO-SLAM [66] introduced global Bundle Adjustment (BA) and loop closure for an implicit SLAM system using an MLP representing a signed-distance field, greatly increasing tracking performance compared to previous radiance field-based SLAM methods. SLAM systems using 3D Gaussian Splatting (3DGS) [21] as their scene representation have also recently been developed [20; 59; 30; 16]. By using 3DGS these systems can achieve higher fidelity scene representations without sacrificing speed. We use MonoGS [30] as the backbone of our active SLAM system.
Active SLAM
In the active SLAM task, the agent must construct a trajectory to explore and map the environment. In order to create an accurate map there should be a trade-off between exploring new regions and reducing the uncertainty of the map and pose estimates [22; 40; 3].
Active SLAM systems are typically divided into three components [27; 37] – candidate goal identification, utility computation, and action planning and execution. Frontier-based exploration (FBE) [58] is a widely used technique for proposing candidate goals [47; 2; 53; 49]. The utility calculated in the second stage of active SLAM algorithms generally seeks to capture uncertainty [27; 37] – specific utility functions are often drawn from either Information Theory (IT) [43] or the Theory of Optimal Experimental Design (TOED) [35]. Metrics from IT are based on entropy, whereas those from TOED are based on covariance. Many TOED-based methods formulate objectives using the Fisher Information Matrix (FIM) because its inverse is the Cramer-Rao lower bound of the covariance matrix and the FIM is generally sparser than the covariance [37]. This approach is often taken by filter-based SLAM algorithms [10; 55; 22], as the filter produces a covariance of the state estimate. For the action planning and execution, various path planning algorithms are used, including RRT [24] in Vallvé & Andrade-Cetto [54]; Huang & Gupta [17] and A [13] in Kim & Eustice [22]. We follow this three-component approach to active SLAM, using standard techniques for the first and third components (FBE and A, respectively) and a novel utility formula based on Fisher Information.
@@
The vast majority of previous work on uncertainty quantification for radiance fields has been for post-processing scenes [44; 45; 50; 12], view selection [34; 19; 50; 25] or active view selection [34; 50; 19], all of which assume the input images are posed. Active neural mapping [61] uses neural variability, that is the prediction robustness against random weight perturbations, as an estimate of uncertainty to actively map a scene with ground truth poses provided. Fisher-RF [19] also performs active scene mapping with ground truth poses provided, based on an approximation of the Fisher Information of views along candidate paths. Zhan et al. [64] perform active reconstruction without ground truth camera poses, however unlike us they only evaluate on small scenes and limit the camera trajectories to be inwards facing. Unlike our method, these prior works only model scene uncertainty, not localization uncertainty which is a key consideration for active SLAM.
@@ 3 Method
Figure 1: An Illustration of Our Active SLAM System The Active Gaussian SLAM system proposes paths based on the Fisher Information about the parameters of the 3D Gaussian parameters. The best path and action along the path is selected with respect to both the information gain and localization accuracy. The Active SLAM system progressively improves the mapping and localization during exploration
In this section, we first introduce the preliminary background of 3D Gaussian Splatting SLAM [30] and FisherRF [19], which are the foundation for our algorithm in Section 3.1. We then discuss our information-driven path proposal algorithm in Sec. 3.2 and our path selection algorithm in Sec. 3.3.
@@ 3.1 Preliminary
In 3D Gaussian Splatting (3DGS) [21], the scene is represented by a set of 3D Gaussians whose color and opacity are learned via a rendering loss. The rendered image depends on the 2D Gaussians $\mathcal{N}(\bm{\mu}_{I},\bm{\Sigma}_{I})$ that are projections $\pi(.)$ of 3D Gaussians $\mathcal{N}(\bm{\mu}_{W},\bm{\Sigma}_{W})$ in world coordinates. The projection of a 3D Gaussian in the world frame to a 2D Gaussian on the image plane can be written as
[…]
$C_{p}=\sum_{i=1}^{N}{\bf c}_{i}\alpha_{i}\prod^{i-1}_{j=1}(1-\alpha_{j}),$ (2)
where ${\bf c}_{i}$ is the learned color associated with each Gaussian, and $\alpha_{i}$ is calculated using the covariance of the corresponding 2D Gaussian following Yifan et al. [62] then multiplying with a learned per-point opacity. Following MonoGS [30], we omit the view-dependence of the color as it greatly saves the number of parameters and memory usage in large-scale scenes while maintaining satisfactory performance. The rendering at a given camera location ${\bf x}$ with 3DGS parameters ${\bf w}$ can be considered as a function $f({\bf x},{\bf w})$.
MonoGS [30] found the Jacobian of the current camera pose ${\bf x}$ with respect to the parameters of 3D Gaussians $\bm{\mu}_{I}$ and $\bm{\Sigma}_{I}$ using the chain rule:
$\frac{\partial{\bm{\mu}_{I}}}{\partial{{\bf x}}}$ $=\frac{\partial{\bm{\mu}_{I}}}{\partial{\bm{\mu}_{C}}}\frac{\mathcal{D}{\bm{\mu}_{C}}}{\mathcal{D}{{\bf x}}}~{},$ (3)
$\frac{\partial{\bm{\Sigma}_{I}}}{\partial{{\bf x}}}$ $=\frac{\partial{\bm{\Sigma}_{I}}}{\partial{\mathbf{J}}}\frac{\partial{\mathbf{J}}}{\partial{\bm{\mu}_{C}}}\frac{\mathcal{D}{\bm{\mu}_{C}}}{\mathcal{D}{{\bf x}}}+\frac{\partial{\bm{\Sigma}_{I}}}{\partial{\mathbf{W}}}\frac{\mathcal{D}{\mathbf{W}}}{\mathcal{D}{{\bf x}}}~{},$ (4)
Let $\tau\in\mathfrak{se}(3)$ and define the (left) partial derivative on the manifold as:
$\frac{\mathcal{D}{f({\bf x})}}{\mathcal{D}{{\bf x}}}\triangleq\lim_{\tau\to 0}\frac{\text{Log}(g(\text{Exp}(\tau)\circ{\bf x})\circ g({\bf x})^{-1})}{\tau}~{},$ (5)
where $\circ$ is a group composition, and Exp and Log are the exponential and logarithmic mappings between Lie algebra and Lie Group.
@@
where $f({\bf x},{\bf w})$ is the rendering model. Under regularity conditions [41], the Fisher Information of $-\log p({\bf y}|{\bf x};{\bf w})$ is the Hessian of Eq. 6 with respect to ${\bf w}$, denoted ${\mathbf{H}}^{\prime\prime}[{\bf y}|{\bf x},{\bf w}]$.
@@
where ${\bf w}$ is the initial estimate of model parameters using current training set $D^{train}$. ${\mathbf{H}}^{\prime\prime}[{\bf w}|D^{train}]^{-1}$ can be computed by summing the Hessians of model parameters across all different views in $\{D_{train}\}$ before inverting. The key of this algorithm is that the Fisher Information ${\mathbf{H}}^{\prime\prime}[{\bf y}^{acq}_{i}|{\bf x}^{acq}_{i},{\bf w}]$ does not depend on the label ${\bf y}^{acq}_{i}$ of the acquisition sample ${\bf x}^{acq}_{i}$. Therefore, it is feasible to compute the Expected Information Gain (EIG) before visiting the potential view candidate ${\bf x}^{acq}_{i}$. However, the number of optimizable parameters is typically more than 20 million, which means it is impossible to compute without sparsification or approximation. In practice, FisherRF [19] applies a Laplace approximation [6; 28] that approximates the Hessian matrix with its diagonal values plus a log-prior regularizer $\lambda I$
@@
To understand the usage of EIG, In Fig. 2, we plot the Peak-signal-to-noise ratio (PSNR) vs EIG at sampled poses and show some example renderings to give a sense of how the EIG is related to the PSNR and the rendering quality. More importantly, unlike PSNR, the EIG can be computed without ground truth images, making it possible to perform view selection during exploration.

### v2

# Multimodal LLM Guided Exploration and Active Mapping using Fisher Information

@@ Abstract
We present an active mapping system that could plan for long-horizon exploration goals and short-term actions with a 3D Gaussian Splatting (3DGS) representation. Existing methods either did not take advantage of recent developments in multimodal Large Language Models (LLM) or did not consider challenges in localization uncertainty, which is critical in embodied agents. We propose employing multimodal LLMs for long-horizon planning in conjunction with detailed motion planning using our information-based algorithm. By leveraging high-quality view synthesis from our 3DGS representation, our method employs a multimodal LLM as a zero-shot planner for long-horizon exploration goals from the semantic perspective. We also introduce an uncertainty-aware path proposal and selection algorithm that balances the dual objectives of maximizing the information gain for the environment while minimizing the cost of localization errors. Experiments conducted on the Gibson and Habitat-Matterport 3D datasets demonstrate state-of-the-art results of the proposed method.
@@ 1 Introduction
Being able to autonomously explore and map an environment while localizing within that map is a core skill for a mobile robot. This ability could empower embodied artificial intelligence systems with effective 3D scene understanding through use in conjunction with open-vocabulary semantic segmentation or vision-language features [91, 68, 55, 87] and can be used as a basis for language-specified robotics tasks [25, 63, 37, 26]. This task is challenging because it requires general knowledge of the typical layout of an environment to identify coarse targets for long-term exploration and the ability to accurately plan each step to ensure localization accuracy while maximizing information gain or ‘surprise’ when traversing the environment.
In this paper, We propose to use 3D Gaussians as our scene representation and tackle this problem in two phases: long-horizon planning with an LLM and detailed motion planning with our information-based approach by quantifying the Fisher Information for 3DGS parameters and current localization states. We prompt a multimodal LLM with our current map state, trajectory, and candidates for long-term exploration frontiers to exploit the LLM’s world knowledge about scene layouts. We then follow the suggestion from the LLM by choosing paths that maximize the Expected Information Gain calculated using the Fisher Information about poses and 3D Gaussians.
Figure 1: An overview about the interaction with a multimodal LLM. The image on the left is a bird’s-eye view rendering from our 3DGS representation. A detailed example of the dialog can be found in the supplementary.
Classical active mapping methods often define objectives to reduce the uncertainty of the state estimate based on its covariance [20, 58, 5, 78, 11, 1]. The covariance is readily available for classical systems which use filters to update the state estimate. However, many recent mapping systems use non-linear optimization to update the state estimate, making the covariance difficult to obtain. In particular, systems [44, 83, 32, 27] with 3DGS [33] for the scene representation have been developed for high-fidelity rendering of novel views of the scene. Previous methods attempted to quantify the uncertainty of radiance fields for reconstructing scenes from given data [66, 67, 72, 23] for active view selection [51, 31, 72, 36] and active reconstruction [84] or mapping [85] of scenes with given localization, and for 3D reconstruction of small scenes with an inward-facing camera [88]. However, all the prior methods only model the uncertainty of the scene representation, ignoring another primary source of error for a practical mapping system: the risk of localization failures. Our proposed system could address not only Information Gain during exploration but also consider localization accuracy when the robot explores unknown and texture-less regions.
In terms of long-horizon planning for robot exploration, classical approaches such as frontier-based exploration and A* algorithms are still used in active mapping systems for their efficiency and simplicity [70, 4, 77, 71, 34]. However, algorithms that use only simple heuristics cannot determine the information gain, nor estimate 3D geometry, limiting their objective to simply improving coverage or having minimal travel distances. To address this issue, learnable approaches have been trained for exploration tasks [22, 59, 38, 46], but these methods are trained for a limited scene distribution and are not natively designed for a 3D Gaussian representation. Recently, LLMs and Vision Language Models (VLM) have demonstrated extraordinary abilities in visual grounding and logical reasoning and have been studied with various robotic tasks [62]. However, these works either focus on other tasks [9, 16, 14], especially tasks with language inputs [26, 74] and navigation [8] or did not take advantage of photo-realistic 3D representation for multimodal LLMs. On the other hand, Neural Radiance Field and 3D Gaussian Splatting have shown the exciting ability to reconstruct accurate 3D scenes just using posed images. Later works [44, 32] enabled online training with localizations of images. However, the 3DGS model reconstructs the scene from a pure optimization perspective without prior knowledge of the scene and only passively works on RGB-D video inputs. In contrast, we can actively select the best paths for exploration to improve the quality of both reconstruction and localization.
We argue that we can leverage the best of both worlds. We can use high-quality bird’s-eye view rendering from a 3D Gaussian Splatting model as the input of the multimodal LLM to help identify long-horizon exploration goals from frontiers in our scene representation. In this way, we could seamlessly integrate multimodal LLM into a path-planning model to provide knowledge of the typical layout of a scene and potential regions of interest for exploration. Besides, we also alleviate the issue by asking the multimodal LLM to provide precise coordinates, which require precise 3D understanding and are dependent on the scale and relative location of the scene. Given the long-horizon exploration goal from the output of the multimodal LLM, we formulate our path planning problem as an active learning problem and propose multiple paths leading to the exploration goal. We consider the expected information gain (EIG) along the planned trajectory with Fisher Information. As shown in Fig. 2, our EIG metric could reflect the ‘surprise’ and correlate well with the rendering quality on candidate views without actually taking a picture at the candidate location. We also derived localization uncertainty on candidate paths with Cramér–Rao bound to balance the dual problem of exploration and localization. Notably, we compute the full Fisher Information matrix for localization parameters without any approximation with our efficient CUDA implementation. Thus, finding the path that most improves scene reconstruction quality when taking observations during exploration while ensuring the agent still has accurate localization. To validate our approach, we evaluate our method on scenes from the Gibson [81] and Habitat-Matterport 3D [56] datasets quantitatively and qualitatively. We show superior reconstruction quality in various metrics compared to several baselines and recent state-of-the-art methods [7, 53, 22, 85, 82]. Our contributions can be summarized as follows:
  • We present an active mapping system that could autonomously explore the environment and extensively compare our system with previous approaches. To the best of our knowledge, we are the first active mapping system with 3D Gaussian representation that is not dependent on ground truth camera pose readings.
  • We provide a way to leverage the zero-shot long-horizon planning ability of LLM into our active mapping system seamlessly.
  • We introduce localization uncertainty with active mapping systems and effectively balance the information gain for exploration and the cost of possible localization errors.
@@ 2 Related work
Active Mapping and Localization
Efficiently exploring an environment in order to map it while being able to localize in that map is a fundamental problem in robotics. There are many methods that address aspects of this problem separately – Simultaneous Localization and Mapping (SLAM) methods [48, 18, 47, 3, 30, 49, 17, 75] address the mapping and localization. Many exploration methods [22, 85, 10] address the exploration and mapping aspects while assuming poses are provided. Active SLAM methods [70, 4, 77, 71] consider both of these problems. The exploration in these systems is usually driven by a measure of uncertainty [39, 54]; specific utility functions are often drawn from either Information Theory (IT) [65], or the Theory of Optimal Experimental Design (TOED) [52]. Recently learning-based approaches have been developed for active mapping and localization, Active Neural SLAM [7] learns policies to drive exploration and estimate the agent pose, more similarly to us NARUTO [21] uses a SLAM backbone to estimate the pose and uses uncertainty to drive exploration. However, they only consider the reconstruction uncertainty, whereas we choose long-horizon planning goals and consider the localization uncertainty as well as the reconstruction uncertainty when planning the best path to the long-horizon goal.
@@
The vast majority of previous work on uncertainty quantification for radiance fields has been for post-processing scenes [66, 67, 72, 23], view selection [51, 31, 72, 36] or active view selection [51, 72, 31], all of which assume the input images are posed. Active neural mapping [85] uses neural variability, that is the prediction robustness against random weight perturbations, as an estimate of uncertainty to actively map a scene with ground truth poses provided. Fisher-RF [31] also performs active scene mapping with ground truth poses provided, based on an approximation of the Fisher Information of views along candidate paths. Zhan et al. [88] perform active reconstruction without ground truth camera poses. However, they only evaluate small-scale scenes and limit the camera trajectories to be inwards facing and only model scene uncertainty, not localization uncertainty, which is a key consideration for active mapping and exploration.
Robot Planning with Foundation Models
LLMs and VLMs have been used widely in robotics as for language-specified tasks, either as high-level planners in conjunction with other methods for low-level control [28, 26, 40, 69, 2, 57, 16, 15], to create intermediate representations which can be planned over [29] or directly outputting actions [90, 41]. Outside of language-driven tasks, VLMs and LLMs have been used in robotics for providing rewards to drive exploration in Reinforcement Learning [76], and for visual localization [45]. More similar to our task, LLMs and VLMs have also been used for goal-driven navigation, such as searching for specific objects [86, 60, 13, 64].
@@ 3 Method
Figure 3: An Illustration of Our Active Mapping System Our method first identifies long-horizon targets with a multimodal LLM by using novel-view synthesis from our 3DGS representation. Our information-based path proposal method then proposes and selects short-term action sequences based on the Fisher Information about the 3D Gaussian parameters. The best path and action along the path is selected with respect to both the information gain and localization accuracy.
We divide the scene exploration task into two phases: (a) long-horizon planning for the coarse direction of exploration that leads to better coverage and understanding (Sec. 3.2), and (b) detailed trajectory proposal (Sec. 3.3) and path selection (Sec. 3.4) that aims to improve 3D reconstruction and localization from a geometrical perspective. The former task relies more on semantic information and prior knowledge about the possible layout of the environment, which is suitable for a generalist model with prior distributions on the scene. The latter task, however, involves detailed motion planning and is better addressed by considering information gain on fine-level geometries. Fig. 3 shows an overview of our method.
@@ 3.1 Preliminary
In 3D Gaussian Splatting (3DGS) [33], the scene is represented by a set of 3D Gaussians whose color and opacity are learned via a rendering loss. A rendered image can be produced by projecting the Gaussians to 2D and using $\alpha$-blending for the $N$ ordered points on the 2D splat that overlaps each pixel. The Jacobian of the current camera pose ${\bf x}$ with respect to the parameters of 3D Gaussians can be computed by defining the (left) partial derivative on the manifold [44]:
$\frac{\mathcal{D}{f({\bf x})}}{\mathcal{D}{{\bf x}}}\triangleq\lim_{\tau\to 0}\frac{\text{Log}(f(\text{Exp}(\tau)\circ{\bf x})\circ f({\bf x})^{-1})}{\tau}~{},$ (1)
where $\tau\in\mathfrak{se}(3)$, $\circ$ is a group composition, and Exp and Log are the exponential and logarithmic mappings between Lie algebra and Lie Group.
@@
where $f({\bf x},{\bf w})$ is the rendering model. Under regularity conditions [61], the Fisher Information of $-\log p({\bf y}|{\bf x};{\bf w})$ is the Hessian of Eq. 2 with respect to ${\bf w}$, denoted ${\mathbf{H}}^{\prime\prime}[{\bf y}|{\bf x},{\bf w}]$. In our formulation ${\bf w}$ is a tensor for all the 3D Gaussian Parameters $R^{N\times 14}$, where $N$ is the number of Gaussians, and ${\bf y}\in R^{\it{h}\times\it{w}\times\it{c}}$ is the RGB-D observation.
@@
where ${\bf w}$ is the initial estimate of model parameters using the current training set $D^{train}$. ${\mathbf{H}}^{\prime\prime}[{\bf w}|D^{train}]^{-1}$ can be computed by summing the Hessians of model parameters across all different views in $\{D_{train}\}$ before inverting. The key of this algorithm is that the Fisher Information ${\mathbf{H}}^{\prime\prime}[{\bf y}^{acq}_{i}|{\bf x}^{acq}_{i},{\bf w}]$ does not depend on the label ${\bf y}^{acq}_{i}$ of the acquisition sample ${\bf x}^{acq}_{i}$. Therefore, it is feasible to compute the Expected Information Gain (EIG) before visiting the potential view candidate ${\bf x}^{acq}_{i}$. However, the number of optimizable parameters is typically more than 20 million, which means it is impossible to compute without sparsification or approximation. In practice, FisherRF [31] applies a Laplace approximation [12, 42] that approximates the Hessian matrix with its diagonal values plus a log-prior regularizer $\lambda I$
@@
3.2 Long Horizon Exploration with Foundation Model
We propose using a multimodal LLM to perform the task of long-horizon planning as a zero-shot planner for the possible paths and leave the detailed path planning for a closed-form uncertainty-aware motion planning algorithm that computes 3D geometries.
As we consider an agent moving in a 2D action plane (e.g., a ground robot), we leverage the view synthesis ability of 3DGS to create expressive bird’s-eye view renderings that could provide an overview of the environment. An occupancy grid on the motion plane of our robot is created using our 3D Gaussian representation, which can be used to identify the frontiers of the current environment. The frontiers are defined as points on the boundary between free space and unobserved space (not marked as free or occupied) in our occupancy map. Each frontier can be defined as a set of neighboring points on the action plane $Q=\{\phi({\bf x}_{i})...\}$, where $\phi(.)$ is the projection from the world frame to the bird’s-eye view. The agent has multiple choices on the frontiers, especially at the early stages of exploration. We apply Chain-of-Thought Prompting [80] to encourage the multimodal LLM to provide analysis on the candidate frontiers first before selection. We provide contextual information such as the description of the task, the total steps allowed for exploration, and our current step in the textual prompt. The multimodal LLM is also allowed to decide whether the robot should not go to a frontier and instead focus on improving existing regions. We embed possible choices of frontiers along with the current location and previously visited trajectory in the bird’s-eye view rendering of the environment along with a text prompt describing the task and contextual information as the input for the multimodal LLM. In this way, we avoid possible infeasible destinations for motion planning. For example, the unexplored regions we have no reconstruction for at this point might be unreachable from the current space. We then parse the long-horizon exploration goal from the reply of LLM. We use GPT-4o [50] for our long-horizon planning task, but our method is agnostic to the underlying multimodal LLM as the inputs for the multimodal LLM are text prompts and a bird’s-eye view rendering. Fig. 1 provides an overview of our interaction with the multimodal LLM. Detailed examples can be found in the Supplementary.